# Neural Operator for Accelerating Coronal Magnetic Field Model

### v2

@@ 4. Conclusion
The exploration of neural operator architectures, particularly the TFNO, represents a significant advancement in the computational modeling of the solar coronal magnetic field. TFNO has demonstrated the capability to simulate the complex magnetic solar outer atmosphere compared to state-of-the-art deep learning methods such as ViT, CNN-RNN, and LSTM. The findings from the comparative study reveal that TFNO achieves superior prediction accuracy as evidenced by the best test performance among all methods. Compared to traditional physics-based models like MHD, TFNO only takes a few seconds to generate the 3D cube while MHD takes days to simulate.
@@
From the physics perspective, TFNO stands out for its precision in predicting the magnetic field topologies, crucially maintaining this accuracy across varying atmospheric heights. TFNO exhibits a remarkable adherence to ground truth, particularly below a height of 2Mm, a region characterized by complex magnetic interactions. Its performance is consistently superior to that of other methods, such as the ViT, CNN-RNN, and CNN-LSTM, especially in critical lower atmospheric layers where precise modeling is most needed. Additionally, TFNO’s alignment with ground truth in the calculation of field orientation metrics further solidifies its capability to capture the nuanced dynamics of solar phenomena.
ACKNOWLEDGEMENT
This work was supported by the National Science Foundation grants AGS-2229064 and AST-2204384, and National Aeronautics and Space Administration Grant 80NSSC20K1282.
Nonlinear Force-Free Fields (NLFFF)
In solar physics, understanding the magnetic field configuration of the solar corona is crucial due to its dominant role in shaping solar phenomena such as flares and coronal mass ejections (CMEs). The coronal magnetic field is often modeled using force-free fields, where the Lorentz force vanishes. This condition implies that the magnetic pressure and tension forces are balanced, making the field force-free.
Mathematically, a force-free magnetic field $\mathbf{B}$ satisfies the equation:
$\mathbf{j}\times\mathbf{B}=0$
where $\mathbf{j}$ is the current density. This condition implies that $\mathbf{j}$ is parallel to $\mathbf{B}$, leading to the formulation:
×$\nabla\times\mathbf{B}=\alpha\mathbf{B}$
Here, $\alpha$ is known as the force-free parameter, which can vary spatially, resulting in a nonlinear force-free field when $\alpha$ is not constant. The variability of $\alpha$ along different field lines is expressed as:
$\mathbf{B}\cdot\nabla\alpha=0$
NLFFF models are essential for accurately reconstructing the solar coronal magnetic field, utilizing boundary conditions obtained from photospheric magnetic field observations. These models are vital for predicting space weather and understanding the mechanisms behind solar eruptions.
The NLFFF database (kusano2020physics), hosted by Nagoya University, provides a comprehensive collection of coronal magnetic field data derived from the Solar Optical Telescope (SOT) onboard the Hinode satellite. This dataset includes three-dimensional magnetic field configurations reconstructed from photospheric vector magnetograms using various NLFFF extrapolation methods.
Figure 8. The distribution of active regions type of training and test split.
[…]
Example.
For a demonstration purpose, we aim to reproduce the $B_{z}$ cube from the boundary condition. We applied TFNO to the active regions from 2010-2014 and used active regions from 2015-2016 as test data. Figure 8 summarizes the distribution of active regions type of training and test split. For the test performance, our TFNO can achieve R-squared 0.9522, Relative Error 0.1752, MSE 0.0477, MAE 0.1018, PSNR 49.05, and SSIM 0.9735.
Figure 9. The visualization of prediction of $B_{z}$ from TFNO compared with ground truth at different heights.
[…]
Example.
Figure 9 shows the visualization of the prediction of $B_{z}$ from TFNO compared with ground truth at different heights. The selected slices are captured at different heights of the cube. The first column in Figure 9 is the input data of the TFNO. The second through fourth columns correspond to heights of approximately 4 Mm, 16 Mm, and 64 Mm, respectively. These slices were chosen to demonstrate the diversity of magnetic field strengths at different heights. The first row represents the “ground truth”. The results of TFNO are almost the same as the ground truth.

### v3

@@ V Conclusion
This paper explores the use of neural operator architectures to accelerate the speed of computational modeling of the solar coronal magnetic field. Specifically, In our proposed GL-FNO, the global branch captures the overall structure and dynamics while the local branch focuses on fine-grained features. This fusion enables our GL-NFO to generalize better across different physical conditions while retaining accuracy in areas requiring detailed resolution. GL-FNO has demonstrated the capability to simulate the complex magnetic solar outer atmosphere compared to state-of-the-art deep learning methods such as FNO, U-NO, U-FNO, ViT, CNN-RNN, and LSTM. The findings from the comparative study reveal that GL-FNO achieves superior prediction accuracy as evidenced by the best test performance among all methods. Compared to traditional physics-based models like MHD, GL-FNO only takes a few seconds to generate the 3D cube while MHD takes days to simulate.
@@
From the physics perspective, GL-FNO stands out for its precision in predicting the magnetic field topologies, crucially maintaining this accuracy across varying atmospheric heights. GL-FNO exhibits a remarkable adherence to ground truth, particularly below a height of 2 Mm, a region characterized by complex magnetic interactions. Its performance is consistently superior to that of other methods, such as the ViT, CNN-RNN, and CNN-LSTM, especially in critical lower atmospheric layers where precise modeling is most needed. Additionally, GL-FNO’s alignment with ground truth in the calculation of field orientation metrics further solidifies its capability to capture the nuanced dynamics of solar phenomena.